What is the first term of a sequence when the eighth term is twice the third term and the sum of the first eight terms is 39?
Answer by Boreal(15235)

a8=2a3
a1+a2+a3+a4+a5+a6+a7+a8=39
(8/2)(a1+a8)=39=4(a1+2a3)
a4=1.2a3
a5=1.4a3
a6=1.6a3
a7=1.8a3
a8=2a3
a3=a3
a2=0.8a3
a1=0.6a3
The sum, 39, equals 10.4 a3
Therefore, a3=3.75
a8=7.5
a1=2.25 ANSWER
sum=4*(a1+a8)=4(9.75)=39